OpenStudy (anonymous):

I need help, If anyone could please help me I'd appreciate it:) I will leave equation below.

OpenStudy (anonymous):

Solve the Equation below: \[4^\left( 3x-5 \right)=256\]

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

It's fine, thanks:)

OpenStudy (anonymous):

I don't expect the answer, I would like to know how to solve this kind of equation.

OpenStudy (anonymous):

Do you know what is the common denominator for 9/16

OpenStudy (anonymous):

Nope

sammixboo (sammixboo):

\(\rm \color{#20bd23}{4^{3x~-~5}~=~256}\) First you gotta take log on both sides of the equation, because we are solving for the exponent \(\rm \color{#20bd23}{log4^{3x~-~5}~=~log256}\) Do you know what to do next?

sammixboo (sammixboo):

If not.. I will get back to you. I have to do something and I am not sure how long it will take

OpenStudy (anonymous):

Yes:)

OpenStudy (anonymous):

No worries, go ahead and do what you have to do. Thank you :)

OpenStudy (anonymous):

rewrite 256 as a power to the base of 4. then, since they have the same base, make the exponents equal to each other then solve.

OpenStudy (jdoe0001):

\(\Large \bf 4^{3x-5}=256\qquad {\color{brown}{ 2^8=256}}\qquad 4^{3x-5}=2^8\qquad x=?\)

OpenStudy (jdoe0001):

hmmm actaullly

OpenStudy (jdoe0001):

\(\Large \bf 4^{3x-5}=256\qquad {\color{brown}{ 4^4=256}}\qquad 4^{3x-5}=4^4\qquad x=?\)

OpenStudy (anonymous):

I don't quite understand what I have to do with \[4^{3x-5}=4^{4}\]

sammixboo (sammixboo):

Okay I am back

sammixboo (sammixboo):

\(\rm \color{#20bd23}{4^4~=~256}\) @jackoo

OpenStudy (anonymous):

okay, so since the bases are the same, 3x-5=4. solve for x

OpenStudy (anonymous):

@sammixboo Okay, I understand that, @iishineontheinside oh okay

sammixboo (sammixboo):

Yes, to solve for x in \(\rm \color{#20bd23}{3x~-~5~=~4}\), you first have add 5 on both sides of the equation \(\rm \color{#20bd23}{3x~-~5~+~5~=~4~+~5}\) \(\rm \color{#20bd23}{3x~=~4~+~5}\) Do you know what to do next?

OpenStudy (anonymous):

Yes, I would have to divide the 3 on both sides giving me x=3

sammixboo (sammixboo):

Correct :)

OpenStudy (anonymous):

That would be all?:)

OpenStudy (anonymous):

that would be all for the question

OpenStudy (anonymous):

I see, thank you all very much:)